### v1

@@ Abstract
Blockchain ecosystems—such as those built around chains, layers, and services—try to engage users for a variety of reasons: user education, growing and protecting their market share, climbing metric-measuring leaderboards with competing systems, demonstrating usage to investors, and identifying worthy recipients for newly created tokens (airdrops). A popular approach is offering user quests: small tasks that can be completed by a user, exposing them to a common task they might want to do in the future, and rewarding them for completion. In this paper, we capture blockchain data from one deployed quest system that offered 43 unique quests over 10 months with 80M completions. We use this data to offer insight about the factors that impact task completion: amount of reward, monetary value of reward, difficulty, and cost. We also discuss the role of farming and bots, and the factors that complicate distinguishing real users from automated scripts.
@@ 1 Introductory Remarks
Blockchain projects understand that even the best technical ideas and code can only achieve so much without user adoption of the platform. Projects struggle to reach users, who generally operate pseudonymously, and users do not always browse for new blockchain projects—there is no high profile ‘DApp store.’ One approach for introducing users to a new blockchain (including independent chains, sidechains, and layer 2 solutions) is showcasing its features and partners through gamification. A set of quests or tasks or boosts are created that offer a bounty or reward to users for performing a specific action on the blockchain, such as transferring a token or using a specific DeFi service. The reward could be a collectable badge (often as an NFT), points on a leaderboard (often referred to as XP, inspired by games like Fortnite and platforms like Duolingo), or an airdrop of tokens with some monetary value. Most projects favour the term ‘quests’ over ‘tasks’ to avoid associations with employment. The term quest highlights the user’s journey toward achievement and mastery, creating a more legally favourable framing rooted in gamification. Despite the aspirational tone, these actions remain structured and transactional, serving both users’ reward-driven motivations and projects’ growth and engagement objectives.
@@
In this measurements paper, we study one quest system over time to understand how the completion-rate of quests are influenced by factors such as difficulty, cost, reward amount, and reward type (monetary or non-monetary value). From our data, we are not able to directly classify users between humans and bots, however we do offer some insights and discussion points about the influence of automated task-completion. To inform this discussion, we identify key stakeholders involved (such as the blockchain, services on the blockchain, investors, activity monitoring services, users, and bots) and analyze their preferences and incentives. We maintain a neutral position on whether quests are useful or not—we study them because they are commonly used (perhaps poised for broader adoption) and we failed to find another detailed study on the parameters that impact task completion.
1.1 Research Questions.
Our paper sheds light on the following research questions (RQs):
  • RQ1: How does the reward amount impact completion?
  • RQ2: How does task difficulty impact completion?
  • RQ3: How does cost to the user impact completion?
  • RQ4: How does enforcing a minimum threshold of value held by the account impact completion?
  • RQ5: How does offering tokens of monetary value impact completion?
  • RQ6: To what extent can the design of the quest help distinguish regular users from automated bots and farming?
@@ 1.2 Related Work
Recent papers have studied airdrops of new tokens [4, 5, 6]. Airdrops studied in this literature are large-scale events where a large quantity (majority to totality) of tokens are distributed to users for the first time. This literature finds a positive impact of airdrops on token value, that retaining user engagement after an airdrop is difficult, and evidence of artificial engagement to ‘farm’ airdropped tokens. If a quest system predates an airdrop, quest completion can be a factor in determining airdrop recipients.
Our research questions are influenced by a persuasion model called the Fogg Behaviour Model [3] which suggests certain deterrents to engagement: time, money, physical effort, brain cycles, social deviance, and non-routine. We examine the influence of quest difficultly (time, brain cycles, and non-routine) and cost (money) on completion, assuming most quests have similar levels of physical effort (minimal) and social deviance (legal or unregulated in most jurisdictions).
1.3 Dataset and Overview
@@
For our paper, we study a quest system built and maintained by a company offering a new blockchain. The quests themselves were co-designed with various services (DApps) running on the blockchain. All of our measurements are replicable from the public website and blockchain data. Our research analyzes data spanning the entirety of 2024. New quests are offered for an epoch, which was initially a one day period and transitioned (at epoch 10) to one week. Our data is organized by epoch 5 (no data was collected for epoch 1–4) to epoch 42. Over this period, 43 unique quests were offered. Some quests were offered in multiple epochs and could be completed multiple times. 1.2M unique addresses completed at least one quest. Across the period, 80M quests were completed.
@@
As seen in Figure 1, overall participation seems to be influenced by various public announcements. When a blockchain does not offer a token but does offer quests, speculative users might assume that an airdrop of tokens is on the blockchain’s future roadmap and quest completion points might factor into who receive airdropped tokens. At epoch 21, the project announces an airdrop and that a snapshot of user activity had been captured as a basis of for the airdrop. Since the snapshot is taken, it is too late to complete quests for the purposes of gaining in airdropped tokens (unless if you speculate that the project will do a second snapshot or a second airdrop). Participation dropped 29.5% immediately after the snapshot announcement. By epoch 23, a further 45.7% drop occurred, totalling a 61.7% decline over two weeks. At epoch 28, a request for proposals (RFP) was announced to ask ecosystem apps and partners to submit how they want to do airdrop allocation, which further cements the fact that quest completions after the snapshot will not be considered. A 22.9% drop in participation follows the RFP. At epoch 30, the RFP closed and the airdrop itself starts, which sees another 42% decline in quest completions. In summary, from epoch 21 to epoch 32, quest completions declined 93.5%.
@@
In epoch 36, a new program was launched where certain quests would receive token rewards (which have monetary value), as opposed to just receiving XP points. We will study this further in RQ5 but for now, we note that this leads an increase in participation: 34% increase in epoch 37 and 128.7 % by epoch 42. The reward program kept evolving and increasing rewards over the following epochs. By epoch 50 the cumulative effect of this token program was 251.8%.
2 Research Questions
2.1 RQ1: Reward Amounts in Points
@@
Figure 2: Reward amount (in non-monetary points) of different quests offered from Epoch 5–42. The main takeaway is that completion appears uncorrelated to points, suggesting other factors impact completion.
@@
Points have long been a key element in gamification systems, primarily used to incentivize specific behaviours or actions, such as educational tasks, objectives in a game, or fostering loyalty through purchases at a business. For quests, points are usually presented in leaderboards—a public, ordered ranking of users and their points, which can boost competition, and result in users striving not merely to collect points but to climb the ranks. Leaderboards have a downside too. Newcomers might feel deterred by the seemingly insurmountable gap between themselves and top-ranked users, and excessive focus on ranking can foster cheating or the transacting of user profiles (with their points).
@@
To what extent do points motivate users? Consider Figure 2 where it shows a dot for each quest offered in each epoch (x-axis) and the number of times it was completed (y-axis). The shape of the curve follows the overall trend (Figure 1). The dots in Figure 2 are colour-coded for how many points the user receives for completing the quest. A simple hypothesis is that quest completions would correlate positively with the number of points. Using the Spearman correlation, we found that correlations across epochs ranged from -0.115 to -0.870, with no epoch showing a significant positive monotonic relationship. In other words, the rank ordering of points and completions does not reveal any consistent patterns—for instance, higher points do not reliably correspond to more completions.
@@
Finally, we note that in the later epochs, a strong negative monotonic relationship emerged, indicating that higher-point quests were completed less frequently. This was an intentional design decision, as more challenging quests were assigned higher point rewards to incentivize completion. However, the data suggests that higher rewards alone were not sufficient to drive engagement for these harder quests. This implies that quest difficulty, along with other factors that act as deterrent, plays a greater role in influencing completions. This finding highlights the need to consider additional factors (RQ2–RQ5) to better predict task completion.
2.2 RQ2: Task Difficulty
@@
Figure 3: The difficulty of a quest is a strong predictor of completion rates, with easy quests (purple) completed most often and difficult quests (orange) completed least often. Variance in the right tail (epochs 40+) correspond to offering medium quests (pink) with tokens of monetary value (see RQ5).
@@
The first research question addresses points, which pull users toward certain quests. On the other side, frictions or deterrents may push users away from certain quests. RQ2–4 address three of these: difficulty, cost, and minimum token requirements (or thresholds). In RQ2, we consider task difficulty. We classify all quests based on (i) completion time, as measured by lead researcher who completed each quest modelling average user behaviour; and (ii) external dependencies that are out of the user’s direct control, such as waiting for a response from another system, interacting with other users, or relying on network availability:
@@
As evident from Figure 3, difficulty is a strong predictor of task completion. Using the Spearman correlation (with reversed difficulty mapping), we found that correlations across epochs ranged from 0.121 to 0.663, with most epochs showing a positive monotonic relationship. As shown in Figure 3, the rank ordering of difficulty and completions reveals consistent patterns—for instance, easier tasks reliably correspond to higher completion rates. This finding reinforces the idea that quest difficulty is a key factor influencing task completion.
However, in certain epochs, the positive correlation between task difficulty and completion rates was not as strong. As shown in Figure 3, some medium or hard quests were completed more frequently than easier tasks, indicating that other deterrent or negative incentives may have a stronger influence on quest completion than difficulty alone. This observation is further explored by analyzing factors such as cost and thresholds in subsequent sections.
2.3 RQ3: Cost
Figure 4: The cost of a quest is a strong predictor of completion rates, with cheap quests (purple) completed most often and costly quests (orange) completed least often.
@@ 2.3 RQ3: Cost
In Figure 4, we show all quests from epochs 37–46 (later quests demonstrated more variety in complexity) and coded their cost into three colours: purple is less than \$0.10 USD, pink is less than \$1.00 USD (and more than purple), and orange is \$1.00 USD and more (no quest was greater than \$10 USD but some cost greater than \$5.00 USD). The analysis of quest completion data revealed a clear trend: quests with a cost under \$0.10 USD are consistently completed by a larger number of users across all epochs, while those costing between \$1 and \$10 USD see minimal engagement. This pattern suggests that cost is a critical deterrent for users in engaging with high-cost quests. As it may lead to a predominantly negative experience for users facing high-fee tasks, quest designers should give consideration to costs.
Using the Spearman correlation, we found that correlations across epochs ranged from -0.843 to -0.896, consistently showing a strong negative monotonic relationship. As shown in Figure 3, the rank ordering of cost and completions reveals a clear pattern—less costly quests (e.g., purple, costing less than \$0.1) are completed far more frequently than higher-cost quests (e.g., orange, costing up to \$10). This finding reinforces the idea that cost is a key deterrent to task completion.
Finally, the data indicates that higher-cost quests significantly deter engagement, suggesting that financial costs plays a critical role in user behaviour and acts as one of the strongest factors influencing completion rates. To encourage participation in higher-cost quests, designers may need to consider additional incentives other than higher point rewards (as seen in the previous section, points alone are not a strong motivator). However, since the relationship between cost and completion is not perfectly linear, further investigation is needed to explore additional financial deterrents—specifically thresholds—to better understand their impact on user engagement and optimize quest design.
2.4 RQ4: Thresholds
Figure 5: The implementation of a minimum threshold of tokens for quest completion has an immediate negative impact on completion when implemented (orange segment of line) on quests that ran without thresholds (purple segment of line). The grey lines show highly completed tasks as a control group.
We consider one last complication for users: quests that can only be completed by accounts with a minimum threshold of tokens. Thresholds are used to combat farming and bots, which we discussed further in the next section. For now, we simply measure their impact.

### v2

@@ Abstract
Blockchain ecosystems—such as those built around chains, layers, and services—try to engage users for a variety of reasons: user education, growing and protecting their market share, climbing metric-measuring leaderboards with competing systems, demonstrating usage to investors, and identifying worthy recipients for newly created tokens (airdrops). A popular approach is offering user quests: small tasks that can be completed by a user, exposing them to a common task they might want to do in the future, and rewarding them for completion. In this paper, we analyze a proprietary dataset from one deployed quest system that offered 43 unique quests over 10 months with 80M completions. We offer insights about the factors that correlate with task completion: amount of reward, monetary value of reward, difficulty, and cost. We also discuss the role of farming and bots, and the factors that complicate distinguishing real users from automated scripts.
@@ 1 Introductory Remarks
Blockchain projects understand that even the best technical ideas and code may struggle to have an impact without user adoption of the platform. There is no clear method for reaching pseudonymous users, who rarely seek out new blockchain platforms, and there’s no prominent ‘DApp store.’
One approach for introducing users to a new blockchain (including layer 1 chains, sidechains, and layer 2 solutions) is showcasing its features and partners through gamification. A set of quests or tasks or boosts are created that offer a bounty or reward to users for performing a specific action on the blockchain, such as transferring a token or using a specific DeFi service. The reward could be a collectible badge (often as an NFT), points on a leaderboard (often referred to as XP, inspired by games like Fortnite and platforms like Duolingo), or an airdrop of tokens with some monetary value [9, 12, 17]. Most projects favour the term ‘quests’ over ‘tasks’ to avoid associations with employment. The term quest highlights the user’s journey toward achievement and mastery, creating a more legally favourable framing rooted in gamification. Despite the aspirational tone, these actions remain structured and transactional, serving both users’ reward-driven motivations and projects’ growth and engagement objectives.
@@
In this measurements paper, we study one quest system over time to understand how the completion-rate of quests is correlated with factors such as difficulty, cost, reward amount, and reward type (monetary or non-monetary value). From our data, we are not able to directly classify users between humans and bots, however we do offer some insights and discussion points about the influence of automated task-completion. To inform this discussion, we identify key stakeholders involved (such as blockchain platforms, blockchain-based services, investors, activity-monitoring services, users, and bots) and analyze their preferences and incentives. We maintain a neutral position on whether quests are useful or not—we study them because they are commonly used (perhaps poised for broader adoption) and we failed to find another detailed study on the parameters that correlate with task completion.
1.1 Research Questions
Our paper sheds light on the following research question (RQ) and two discussion questions (DQs). DQs have less (or no) empirical basis and are instead based on insights or discussion points.
  • RQ1: What factors are correlated with quest completion rates? We analyze (a) reward amount, (b) task difficulty, (c) cost, (d) reserve requirements, and (e) rewards with monetary value.
  • DQ1: To what extent can the design of the quest help distinguish genuine users from automated bots and farming?
  • DQ2: Who are the stakeholders in quest systems? What can we learn from their preferences?
@@ 1.2 Related Work
We are not aware of other research papers on quest systems for blockchain projects. Brand loyalty programs [6] have been studied deeply, and more recently gamification of learning tasks [10]. Experience points in games have been criticized [3]. For loyalty programs, Wulf et al. outline several drivers of participation [16]. Relevant to our study are: users engage when costs are zero as opposed to non-zero, they are asked to make minimal efforts as opposed to extended efforts, and when rewards are ‘hard’ (e.g., a discount) as opposed to ‘soft’ (e.g., product updates).
@@
On the blockchain side, the relationship between new tokens and user adoption has been studied [2]. One particular (and recent) focus is on airdrops [9, 12, 17] where a large quantity (majority to totality) of tokens are distributed to users for the first time. This literature finds a positive impact of airdrops on token value, that retaining user engagement after an airdrop is difficult, and evidence of artificial engagement to ‘farm’ airdropped tokens (including detection algorithms [7, 18]). In our dataset, the quest system is run both before and after an airdrop. In discussing the role of farming and bots, we use a persuasion model [8] which suggests certain deterrents to engagement: time, money, physical effort, brain cycles, social deviance, and non-routine. We examine the relationship between quest difficulty (time, brain cycles, and non-routine) and cost (money) with completion, assuming most quests have similar levels of physical effort (minimal) and social deviance (legal or unregulated in most jurisdictions).
1.3 Dataset & Limitations
Reproducibility.
For our paper, we study a quest system built and maintained by a company offering a new EVM-compatible, proof of stake blockchain. The quests themselves were co-designed with various services (DApps) running on the company’s blockchain. The choice to use proprietary data has pros and cons.
The disadvantage is that we offered the company final say over being named or identifiable in our paper; they declined after our study was complete. Thus we only provide our dataset after processing and de-identifying. [^1] Available on GitHub: https://github.com/MadibaGroup/2025-Quests Our statistics and figures can be replicable from the data and new insights could be developed by other researchers, however the accuracy of the data cannot be independently verified against on-chain data or website archives.
The advantage of engaging with a company is that we were able to inform the quest design to some extent. For example, we were able to suggest the addition of reserve requirements (§2.4) to quests and rewards with monetary value (§2.5), and ensure they were added/removed from tasks that otherwise were unchanged (isolating two main variables: the requested feature and the epoch). To control for different completion rates across epochs, we also ensured some tasks were offered across multiple epochs without change.
Representativeness.
Our dataset originates from a single platform, which means we cannot assert broad generalizability. However, the platform’s quests and bot-control measures appear standard and comparable to other blockchain quest systems we examined. Our data covers both pre- and post-airdrop periods, enabling us to observe user behaviour changes around significant events. It’s important to note that our findings may become less representative over time due to the increasing ease with which non-experts can create bots, particularly as AI-driven agents [1, 15] become more prevalent.
Ground Truth on Bots.
We lack ground truth distinguishing genuine users from Sybil accounts, whether manual (farming) or automated (bots/scripts). Thus, our data reports only unique address completions, and interpretations should be made cautiously of this fact. Nevertheless, understanding factors correlated with quest completion rates remains valuable for designing effective blockchain engagement strategies, even without explicit bot classification.
1.4 Data Analysis
The data we used includes the quest system’s website interface where each quest corresponds to specific actions executed on-chain, defined by interactions with smart contracts. Upon quest completion, the loyalty program verifies the user’s on-chain transaction to confirm that all quest requirements are met. Once verified, the program stores the details in the company’s database—including transaction hash, user address, and contract address. This information is merged with additional off-chain data such as the points awarded for quest completion and the epoch, a time metric controlled by the loyalty program rather than the blockchain. We were provided SQL access to this raw data and exported CSV files (anonymized and available on our repository) for analysis. The statistics we compute (such as Spearman correlation) are documented in our repository. We skip the details in the paper as the techniques themselves not novel or innovative.
1.5 Key Events in Dataset
@@
Our research analyzes data spanning the entirety of 2024. New quests are offered for an epoch, which was initially a one day period and transitioned (at epoch 10) to one week. Our data is organized by epoch 5–42 (the company’s backend database starts at epoch 5). Over this period, 43 unique quests were offered. Some quests were offered in multiple epochs and could be completed multiple times. 1.2M unique addresses completed at least one quest. Across the period, 80M quests were completed.
@@
As seen in Figure 1, overall participation seems to be influenced by various public announcements. When a blockchain does not offer a token but does offer quests, speculative users might assume that an airdrop of tokens is on the blockchain’s future roadmap and quest completion points might factor into who receives airdropped tokens. At epoch 21, the project announces an airdrop and that a snapshot of user activity had been captured as a basis for the airdrop. Since the snapshot is taken, it is too late to complete quests for the purposes of gaining in airdropped tokens (unless you speculate that the project will do a second snapshot or a second airdrop). Participation dropped 29.5% immediately after the snapshot announcement. By epoch 23, a further 45.7% drop occurred, totalling a 61.7% decline over two weeks. At epoch 28, a request for proposals (RFP) was announced to ask ecosystem apps and partners to submit how they want to do airdrop allocation, which further cements the fact that quest completions after the snapshot will not be considered. A 22.9% drop in participation follows the RFP. At epoch 30, the RFP closed and the airdrop itself starts, which sees another 42% decline in quest completions. In summary, from epoch 21 to epoch 32, quest completions declined 93.5%.
@@
In epoch 36, a new program was launched where certain quests would receive token rewards (which have monetary value), as opposed to just receiving XP points. We will study this further later but for now, we note that this leads an increase in participation: 34% increase in epoch 37 and 128.7 % by epoch 42. The reward program kept evolving and increasing rewards over the following epochs. By epoch 50 the cumulative effect of this token program was 251.8%.
Retention rate can be computed as $RR=\frac{(E-N)}{S}\cdot 100$ for users at the start ($S$), end ($E$), and newly acquired users ($N$). Between epochs 21 and 35, RR was -0.183%. After token rewards were introduced (epoch 42), $RR$ rose slightly to just 0.59%. Transaction volume dropped sharply from 11,301,078 at peak participation to just 572,595 by epoch 39 (5% of peak volume). Taken together, retention is extremely low and incentive-driven (reacting to monetary rewards). This supports the idea that user engagement was transactional and profit-driven rather than loyalty-driven.
2 RQ1: Task Completion Factors
2.1 Reward Amounts in Points
@@
Figure 2: Reward amount (in non-monetary points) of different quests offered from Epoch 5–42. Completion rates do not correlate with points, indicating other contributing factors.
@@
Points have long been a key element in gamification systems [10], primarily used to incentivize specific behaviours or actions, such as educational tasks, objectives in a game, or fostering loyalty through purchases at a business. For quests, points are usually presented in leaderboards—a public, ordered ranking of users and their points, which can boost competition, and result in users striving not merely to collect points but to climb the ranks. Leaderboards have a downside too. Newcomers might feel deterred by the seemingly insurmountable gap between themselves and top-ranked users, and excessive focus on ranking can foster cheating or the transacting of user profiles (with their points).
@@
To what extent do points motivate users? Consider Figure 2 where it shows a dot for each quest offered in each epoch (x-axis) and the number of times it was completed (y-axis). The shape of the curve follows the overall trend (Figure 1). The dots in Figure 2 are colour-coded for how many points the user receives for completing the quest. A simple hypothesis is that quest completions would correlate positively with the number of points. Using the Spearman correlation [14], we found that correlations across epochs ranged from -0.115 to -0.870, with no epoch showing a significant positive monotonic relationship. In other words, the rank ordering of points and completions does not reveal any consistent patterns—for instance, higher points do not reliably correspond to more completions.
@@
Finally, we note that in the later epochs, a strong negative monotonic relationship emerged, indicating that higher-point quests were completed less frequently. This was an intentional design decision, as more challenging quests were assigned higher point rewards to incentivize completion. As the data suggests, higher point rewards alone were not associated with increased engagement for these harder quests. Instead, quest difficulty, and other factors that act as deterrents, are more strongly associated with variations in task completion. This finding highlights the need to consider additional factors to better understand the predictors of task completion.
2.2 Task Difficulty
@@
Figure 3: Quest difficulty is strongly associated with completion rates, with easy quests (purple) completed most often and difficult quests (orange) completed least often. Variance in the right tail (epochs 40+) correspond to offering medium quests (pink) with tokens of monetary value.
@@
The first research question addresses points, which pull users toward certain quests. On the other side, frictions or deterrents may push users away from certain quests. We willl address three of these: difficulty, cost, and minimum token requirements (or thresholds), starting with task difficulty. We classify all quests based on (i) completion time and (ii) external dependencies that are out of the user’s direct control, such as waiting for a response from another system, interacting with other users, or relying on network availability. Completion time was measured by lead researcher who completed each quest modelling average user behaviour; we felt this was objective enough but more robust followup studies might take the agreement of several researchers independently assessing difficulty.
@@
As evident from Figure 3, difficulty is strongly associated with task completion. Using Spearman correlation (with reversed difficulty mapping), we found that correlations across epochs ranged from 0.121 to 0.663, mostly indicating a positive monotonic relationship. Easier tasks consistently had higher completion rates.
Nevertheless, in certain epochs, the positive correlation was not as strong. As shown in Figure 3, some medium or hard quests were completed more frequently than easier tasks, suggesting that other deterrents or negative incentives may be more strongly associated with quest completion than difficulty alone. This observation is further explored by analyzing factors such as cost and thresholds in subsequent sections.
2.3 Cost
Figure 4: Quest cost is strongly associated with completion rates, with cheaper quests (purple) completed most often and more expensive quests (orange) completed least often.
@@ 2.3 Cost
In Figure 4, we show a selection of quests from epochs 37–46 (later quests demonstrated more variety in complexity) and coded their cost into three colours: purple is less than \$0.10 USD, pink is less than \$1.00 USD (and more than purple), and orange is \$1.00 USD and more (no quest was greater than \$10 USD but some cost greater than \$5.00 USD). Quest costs in USD were computed using token exchange rates at the time of each transaction, based on price data provided by CoinGecko, aligning precisely with the prices displayed to users on the quest platform UI. The analysis of quest completion data revealed a clear trend: quests with a cost under \$0.10 USD are consistently completed by a larger number of users across all epochs, while those costing between \$1 and \$10 USD see minimal engagement. This pattern suggests that cost is strongly associated with lower engagement in high-cost quests. Given this observed relationship, quest designers may benefit from considering alternative incentives or cost structures.
Using the Spearman correlation, we found that correlations across epochs ranged from -0.843 to -0.896, consistently showing the strongest monotonic relationship among all research questions. As shown in Figure 3, the rank ordering of cost and completions reveals a clear pattern—less costly quests (e.g., purple, costing less than \$0.10) are completed far more frequently than higher-cost quests (e.g., orange, costing up to \$10). This finding reinforces the idea that cost is strongly associated with lower task completion.
Since cost showed the strongest correlation, we infer that users are not completing quests randomly, nor are they primarily motivated by mastery or loyalty. Instead, their behavior suggests a tendency to minimize cost, aligning with patterns commonly associated with Sybils or automated bots. This observation led us to the next research question, where thresholds were introduced to user wallets to investigate their association with completion rates and assess whether financial deterrents influence engagement in a more structured manner.
2.4 Thresholds
Figure 5: The implementation of a minimum threshold of tokens for quest completion is associated with an immediate decline in completions for quests that previously ran without thresholds (orange segment compared to the purple segment). The grey lines represent highly completed tasks as a control group.
We consider one last complication for users: quests that can only be completed by accounts with a minimum threshold of tokens. Thresholds are used to combat farming and bots, which we discussed further in the next section. For now, we simply measure their correlation with quest completion.